OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83? -200 -197 -194 -191

OpenStudy (anonymous):

@ghazi @ghazi @ghazi

OpenStudy (ghazi):

ust tn = a+(n-1)d

OpenStudy (ghazi):

*use

OpenStudy (anonymous):

an=a1+(n-1)d since a17 is given it is given that -83=13+16d -96=16d d=-6 a35=a1+34d put value of d

OpenStudy (anonymous):

how would you put the a into the calculator so it goes lower like a reverse to the power of

OpenStudy (anonymous):

i know its like some form of the parenthesis

Directrix (directrix):

Let a1 and a17 and a35 represent terms a sub 1, a sub 17, and a sub 35, respectively.

Directrix (directrix):

The nth term of an arithmetic sequence is the first term + (n-1)*d where d is the common difference of the sequence.

Directrix (directrix):

Terms a1 and a17 are 16 places apart. So, a17 = a1 + 16 d -83 = 13 + 16d -96 = 16d -6 = d Please check this.

OpenStudy (anonymous):

I get that I believe

Directrix (directrix):

an = a1 + (n-1)*d a35 = 13 + (35-1)*(-6)

Directrix (directrix):

a35 = 13 + (35 - 1) *(-6) a35 = 13 + 34 * (-6) = ? --->@tinaballerina

OpenStudy (anonymous):

so -191 , right?

Directrix (directrix):

I got -203. Let me check my sign-us-ses

Directrix (directrix):

13 + 34*(-6) = 13 - 204 = -191

Directrix (directrix):

I multiplied (-6) by 36 for some reason and messed up.

OpenStudy (anonymous):

Ah I see. Glad to see someone takes time to help! Thank you I very much appreciate it, would you mind checking over my few problems? I answered them.

Directrix (directrix):

Are these other problems in separate threads or ones you have not yet posted or what?

OpenStudy (anonymous):

I have not posted them. I have the whole problems I can post them in the comment

Directrix (directrix):

One question per thread works better for me because I have to scroll up and down to get the details of a problem. Then, if several problems and the details thereof intermingle, it makes for a mess.

OpenStudy (anonymous):

Okay Ill link you per problem?

Directrix (directrix):

This @ping system is not working. The send a message via the envelope is working. You can send a message, or I can find you.

OpenStudy (anonymous):

Okay theres not many questions being asked so ill ask in the math lol